# Emergent Turbulence in Nonlinear Gravity

### v1

@@ Discussion
We have unraveled nonlinear phenomena in General Relativity in 3+1 settings. By stirring asymptotically flat spacetimes (with vanishing cosmological constant) through quasi-steady gravitational waves, we identified and quantified the crucial role of three- and four-mode interactions in inducing resonant and anti-damping instabilities, which preferentially channel energy inversely toward larger scales. Our analysis demonstrates that these instabilities are key drivers of emergent gravitational turbulence: the inverse energy cascade in both temporal and angular directions, together with the resonant and anti-damping instabilities revealed here, are essentially two sides of the same coin. Similar mechanisms, and their underlying mathematical structure, are known in turbulent two-dimensional fluids under steady forcing, which, in turn, shed light on decaying turbulence (non-driven) scenarios [3]. While in fluid dynamics the inverse cascade is understood through enstrophy, its gravitational counterpart remains to be fully characterized.
Beyond advancing the fundamental understanding of nonlinear dynamics in General Relativity, our observations are particularly compelling in the context of binary black hole mergers. For instance, an equal-mass, quasi-circular system is predominantly driven by an $\ell=2$ mode whose frequency increases as the orbit tightens due to the loss of energy and angular momentum through gravitational radiation. Throughout the merger, neither theory nor observation shows significant generation of higher $\ell$ modes or higher frequencies (see, e.g. [24, 25]). Such behavior aligns naturally with our results, as the system’s response to the driving tends to preserve the lowest angular structure and frequency.
More broadly, our observations can inform the study of nonlinear gravitational wave scattering. For instance, gravitational scattering of particles has recently led to promising developments in the Post-Minkowskian framework [26], which has become an important tool for understanding the relativistic two-body problem and generating accurate gravitational waveforms. Nonlinear wave scattering, however, has been far less explored, yet may contain rich information relevant to Post-Minkowskian expansions of General Relativity. Also, possible impacts of gravitational turbulence in the early universe have been raised in [27]. We are therefore at the dawn of a new era in gravitational physics, where nonlinearities are poised to play a central role.
@@
In the context of the fluid-gravity correspondence, the turbulent behavior unraveled here clearly extends beyond the established classical fluid-gravity duality scenarios, which rely on the presence of a black hole, a negative cosmological constant, and long-wavelength perturbations relative to the black hole scale [10]. Furthermore, our discovery comprehensibly broadens and deepens earlier hints — based on black hole perturbative studies — of such behavior in highly specialized settings without a cosmological constant [8]. The newly gained understanding, which does not require a black hole or a negative cosmological constant, offers a new platform for studying and comparing nonlinear mechanisms in gravity and fluid mechanics. It may also inspire new directions for exploring emerging dualities, such as celestial holography [28] and connections between Carrollian hydrodynamics and gravity [29].
@@ Acknowledgements.
We thank Luciano Combi for discussions during the initial stage of this work. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work makes use of the Black Hole Perturbation Toolkit. LL acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant. LL also thanks financial support via the Carlo Fidani Rainer Weiss Chair at Perimeter Institute and CIFAR. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
Supplementary Information
Here we provide more information to help understand the results in the main text.
I Methods
@@
$P_{\mu\nu}^{{\rm P}\alpha\beta}d_{t}u^{1-}_{\alpha\beta}\doteq\left({\rm incoming\leavevmode\nobreak\ gravitational\leavevmode\nobreak\ wave}\right)_{\mu\nu},$ (6)
@@
II frequency-dependent mass accretion
This section investigates the time evolution of the black hole’s Christodoulou mass, $M$, under several representative scenarios. In Fig. 3a, we inject gravitational waves with $\ell=2$ (dashed curves) and $\ell=6$ (solid curves) into a Schwarzschild black hole. For $\ell=2$, we consider two frequencies: $\omega=0.5/M_{i}$ (dark blue) and $0.3/M_{i}$ (light blue); whereas for $\ell=6$, the injected frequencies are $\omega=0.5/M_{i}$ (light brown) and $1.5/M_{i}$ (dark brown). Here $M_{i}$ denotes the black hole’s initial mass, so all frequencies have fixed values regardless black hole’s growth.
The gravitational-wave drivers are ramped up gradually at early times (gray region), after which the amplitudes are held constant. Three distinct mass accretion behaviors are observed: (1) For $(\ell=2,\omega=0.5/M_{i})$ and $(\ell=6,\omega=1.5/M_{i})$, the black hole mass grows linearly with time, indicating a steady accretion regime — albeit with different rates. (2) For $(\ell=6,\omega=0.5/M_{i})$, the mass remains constant, signaling no accretion. (3) For $(\ell=2,\omega=0.3/M_{i})$, the growth rate is initially low and time-dependent, but the system eventually transitions into a steady accretion phase at late times.
Figure 3: Mass accretion induced by gravitational-wave drivers. a. Time evolution of the black hole’s Christodoulou mass $M$, normalized by its initial value $M_{i}$, for four representative cases: $\ell=2,\omega=0.5/M_{i}$ (dark-blue dashed), $\ell=2,\omega=0.3/M_{i}$ (light-blue dashed), $\ell=6,\omega=0.5/M_{i}$ (light-brown solid), and $\ell=6,\omega=1.5/M_{i}$ (dark-brown solid). b. Transmissivity of a Schwarzschild black hole for $\ell=2$ (gray) and $\ell=6$ (black), computed using Black Hole Perturbation Toolkit [37]. Arrows indicate the four drivers used in a. c. Accretion rate induced by the $\ell=2,\omega=0.5/M_{i}$ driver as a function of injection amplitude $A_{i}$, which shows a clear quadratic dependence.
In scenario (1), both waves have relatively high frequencies, as indicated by the dark-blue dashed arrow and the light-brown solid arrow in Fig. 3b. Their transmission coefficients are close to unity, allowing nearly all of the wave energy to be absorbed. The black hole thus accretes efficiently and steadily, with its mass increasing linearly in time. In scenario (2), shown by the dark-brown solid arrow, the frequency $\omega=0.5/M_{i}$ is much lower than the quasinormal mode frequency for $\ell=6$. The transmission coefficient is near zero, and the wave is almost entirely reflected — leading to negligible mass growth. Scenario (3) corresponds to the light-blue dashed arrow, where the transmissivity is initially small $(\lesssim 0.3)$ but non-negligible. Only a fraction of the wave is absorbed at early times, resulting in a slow, time-dependent accretion rate. However, as the black hole mass $M$ increases, so does the product $\omega M$, the system moves to the right along the transmission curve [^3] We emphasize that $M$ here refers to the instantaneous Christodoulou mass of the black hole, which evolves during the simulation. This should be distinguished from the fixed initial mass $M_{i}$. The injected frequency $\omega M_{i}$ remains constant throughout.. Eventually, $|T_{\ell}|$ asymptotes to unity, and nearly all the wave energy is absorbed. At this stage, the system transitions to a steady accretion regime with a constant mass growth rate.
In addition to mass growth, we find that the black hole’s spin remains approximately zero $(\sim 10^{-13})$ throughout simulations, as expected for injections of purely gravitoelectric waves. Taken together, these results provide a strong internal consistency check for our numerical setup. They demonstrate that the black hole absorbs gravitational waves and exhibits frequency-dependent accretion behavior in excellent agreement with expectations from linear perturbation theory. When the transmission coefficient approaches unity, the black hole enters a steady accretion regime.
III Case study: $\ell=2$ driver
In the following, we elaborate on these phenomena by presenting additional results and interpretations for both the weak regime (Sec. III.1) and the strong regime (Sec. III.2). In particular, the discussion of the weak regime helps build intuition about our setup and draw on insights from black hole linear perturbation theory and tidal phenomena.
III.1 Weak regime
$\underbrace{(\omega,\ell)\otimes\ldots(\omega,\ell)}_{n}.$ (16)
In our case, they lead to:
  • Linear order: $\omega$ in $\ell=2$, with amplitude $\propto A_{i}$.
$|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}.$ (17)
To our knowledge, Eq. (17) has not been rigorously extended to higher orders. However, it is natural and reasonable to postulate that a $n$-th order effect populates multiple harmonics up to $\ell_{\max}=2n$. A direct consequence is that the $2\omega$ component in $\ell=6$ (see discussions below in Sec. III.3) cannot arise at quadratic order $(2\omega=\omega+\omega)$, but must instead originate from a fourth-order process: $2\omega=\omega+\omega+\omega-\omega$.
III.2 Strong regime
Figure 5 complements our discussions in the main text. Similar to the $\ell=6$ response, where $2\omega$ and $\omega$ emerge sequentially at late times, Fig. 5a shows that in $\ell=4$, the $\omega$ component is nonlinearly excited (orange region) after the laminar modes $2\omega$ and $3\omega$ (gray regions) have peaked and the driver has stabilized. As discussed in Sec. III.1 and Fig. 4, these two laminar modes arise from second- and third-order couplings, respectively. The yellow dashed curve in the bottom panel of Fig. 5a connects the dominant laminar mode $2\omega$ with the later-emerging mode $\omega$, highlighting an inverse cascade.
Within the time interval $0,700M_{i}]$, the amplitude distributions are qualitatively consistent with expectations from the normal perturbation perspective. As discussed in Sec. III.1, in $\ell=2$, each frequency multiple $n\omega$ ($n\in{\rm integer}$) arises from a $n$-th order coupling, whose amplitude decreases exponentially with $n$. In $\ell=4$ (Fig. 6b), the dominant frequency component is $2\omega$, originating at second order. The $\omega$ and $3\omega$ components both appear at third order, and thus have comparable amplitudes, together forming a “ v ” shape. Higher multiples arise from progressively higher-order interactions and show correspondingly smaller amplitudes. Finally, in $\ell=6$ (Fig. 6c), the leading third-order modes are $\omega$ and $3\omega$, followed by two fourth-order modes $2\omega$ and $4\omega$. Modes arising from the same perturbative order have comparable amplitudes, producing a “ N ” shape in the spectrum.
III.3 Injection at various frequencies
Our investigation in Sec. III.2 focused on a specific injection frequency $\omega M_{i}=0.5$, chosen somewhat arbitrarily but intentionally close to and slightly above the $\ell=2$ fundamental quasinormal mode frequency. One motivation for this choice was to ensure that a large portion of the injected wave reaches the horizon. In what follows, we demonstrate that the key features of the emergent instability persist for other low-frequency injections.
Figure 8 shows the amplitude envelopes of several modes undergoing the emergent instability, with injection frequencies $\omega M_{i}$ at 0.4 (a) and $0.6$ (b). We first observe that the $2\omega$ component in $\ell=6$ undergoes a laminar excitation during the initial injection phase ($t<t_{0}=600M_{i}$, in gray). As discussed earlier in Sec. III.1, it is a fourth-order effect: $2\omega(=\omega+\omega+\omega-\omega)$. The peak amplitude of this mode has been confirmed to scale as $A_{\rm inj}^{4}$. After the injection stabilizes $(t>t_{0}=600M_{i})$, $2\omega$ decays to some extent and undergoes the secondary excitation (open circle). The mode grows more rapidly for $\omega M_{i}=0.6$ than 0.4, and in the former case, the emergent excitation dominates over the initial laminar response. Ultimately, the evolution curve of $2\omega$ forms a two-peak structure.
IV Case study: $\ell=6$ driver
  • Linear order: $\omega$ in $\ell=6$, with amplitude $\propto A_{i}$.

### v2

@@ Discussion
We have unraveled nonlinear phenomena in GR in 3+1 settings. By stirring asymptotically flat spacetimes through quasi-steady GWs, we identified the crucial role of three- and four-mode interactions, arising from the full nonlinear structure of Einstein’s equations, in inducing resonant and anti-damping instabilities, which preferentially channel energy inversely toward larger scales. Our analysis demonstrates that the two instabilities are key drivers of emergent gravitational turbulence: the inverse energy cascade in both temporal and angular directions, together with the resonant and anti-damping instabilities revealed here, are essentially two sides of the same coin. Similar mechanisms, and their underlying mathematical structure, are known in turbulent two-dimensional fluids in both driven and decaying turbulence (non-driven) scenarios [25]. While in fluid dynamics the inverse cascade is understood through enstrophy, its gravitational counterpart remains to be fully characterized.
Our findings could offer a new perspective on BBH dynamics: The pre-merger stage acts as a driven stage, while the merger initiates a non-driven/decaying phase that relaxes toward a laminar regime. Just as driven turbulence in fluids provides key insights into its decaying counterpart, our driven setup enables understanding how nonlinear gravitational dynamics settle into linear behavior. Especially, we find that the response to external driving favors the lowest angular structures and frequencies, consistent with theoretical and observational results for BBHs. Future work could drive a BH into a quasi-steady state and then remove the forcing, thereby directly probing the gravitational analogue of the transition from driven to decaying turbulence.
As discussed, our driven setup can establish an “inertial-type” regime and identify key nonlinear mechanisms that are expected to also operate in non-driven cases. This opens a new approach to studying nonlinear dynamics in GR. Meanwhile, the setup mimics astrophysical triple systems [7, 38], where a BH continuously receives GWs from a nearby binary. Our observed turbulence suggests that similar phenomena could arise in such environments. Furthermore, the turbulence may connect to hidden integrability structures in GR [20, 21], nonlinear dynamics in the early universe [14, 15], and nonlinear GW scattering in the post-Minkowskian framework, see e.g., [13].
@@
Within the fluid-gravity correspondence, the unraveled turbulence extends beyond the classical duality, which needs a BH, a negative cosmological constant, and long-wavelength perturbations relative to the BH scale [3]. Our results broaden and deepen earlier perturbative hints of such duality in specialized asymptotically flat settings [44]. This new understanding, which does not require a BH or cosmological constant, offers a new platform for studying and comparing nonlinear mechanisms in gravity and fluid mechanics. It may also inspire explorations of emerging dualities such as celestial holography [35] and the gravity-Carrollian hydrodynamics correspondence [12].
Although our steady gravitational driving and inverse cascades mark a concrete step toward the fluid–gravity correspondence, a fundamental distinction remains between them. Fluid nonlinearities arise from the principal (advective) terms, which allow characteristic crossings in the inviscid case; whereas in GR such crossings never occur, nonlinearity emerges from non-principal terms. In essence, gravity “gravitates,” focusing energy and mediating its nonlinear dynamics. Understanding how this distinction manifests within the correspondence is an intriguing direction.
@@ Acknowledgements.
We thank Luciano Combi for discussions during the initial stage of this work. This work makes use of the Black Hole Perturbation Toolkit. LL acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant. LL also thanks financial support via the Carlo Fidani Rainer Weiss Chair at Perimeter Institute and CIFAR. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
Supplementary Material
I frequency-dependent mass accretion
As shown in Fig. 6a, we investigate the time evolution of the black hole’s Christodoulou mass, $M$, under several representative scenarios. In each case, the driving gravitational wave has a fixed frequency, expressed in units of the black hole’s initial mass $M_{i}$. The driver’s amplitude is gradually ramped up at early times (gray region) and then held constant thereafter.
Three distinct mass accretion behaviors are observed:
  1. 1.
    For $(\ell=2,\omega=0.5/M_{i})$ and $(\ell=6,\omega=1.5/M_{i})$ — shown as the dark-blue dashed and light-brown solid curves — the black hole mass grows linearly with time, indicating a steady accretion regime — albeit with different rates.
  2. 2.
    For $(\ell=6,\omega=0.5/M_{i})$ — dark-brown solid curve — the mass remains constant, signaling no accretion.
  3. 3.
    For $(\ell=2,\omega=0.3/M_{i})$ — light-blue dashed curve — the growth rate is initially low and time-dependent, but the system eventually transitions into a steady accretion phase at late times.
Figure 6: Mass accretion induced by gravitational-wave drivers. a. Time evolution of the black hole’s Christodoulou mass $M$, normalized by its initial value $M_{i}$, for four representative cases: $\ell=2,\omega=0.5/M_{i}$ (dark-blue dashed), $\ell=2,\omega=0.3/M_{i}$ (light-blue dashed), $\ell=6,\omega=0.5/M_{i}$ (dark-brown solid), and $\ell=6,\omega=1.5/M_{i}$ (light-brown solid). b. Transmissivity of a Schwarzschild black hole for $\ell=2$ (gray) and $\ell=6$ (black), computed using Black Hole Perturbation Toolkit [4]. Arrows indicate the four drivers used in a. c. Accretion rate induced by the $\ell=2,\omega=0.5/M_{i}$ driver as a function of injection amplitude $A_{i}$, which shows a clear quadratic dependence.
In scenario (1), both waves have relatively high frequencies, as indicated by the dark-blue dashed arrow and the light-brown solid arrow in Fig. 6b. Their transmission coefficients are close to unity, allowing nearly all of the wave energy to be absorbed. The black hole thus accretes efficiently and steadily, with its mass increasing linearly in time. In scenario (2), shown by the dark-brown solid arrow in Fig. 6b, the frequency $\omega=0.5/M_{i}$ is much lower than the quasinormal mode frequency for $\ell=6$. The transmission coefficient is near zero, and the wave is almost entirely reflected — leading to negligible mass growth. Scenario (3) corresponds to the light-blue dashed arrow in Fig. 6b, where the transmissivity is initially small $(\lesssim 0.3)$ but non-negligible. Only a fraction of the wave is absorbed at early times, resulting in a slow, time-dependent accretion rate. However, as the black hole mass $M$ increases, so does the product $\omega M$, the system moves to the right along the transmission curve [^4] We emphasize that $M$ here refers to the instantaneous Christodoulou mass of the black hole, which evolves during the simulation. This should be distinguished from the fixed initial mass $M_{i}$. The injected frequency $\omega M_{i}$ remains constant throughout.. Eventually, $|T_{\ell}|$ asymptotes to unity, and nearly all the wave energy is absorbed. At this stage, the system transitions to a steady accretion regime with a constant mass growth rate.
In addition to mass growth, we find that the black hole’s spin remains approximately zero $(\sim 10^{-13})$ throughout simulations, as expected for injections of $(\ell=2,m=0)$ waves. Taken together, these results provide a strong internal consistency check for our numerical setup. They demonstrate that the black hole absorbs gravitational waves and exhibits frequency-dependent accretion behavior in excellent agreement with expectations from linear perturbation theory. When the transmission coefficient approaches unity, the black hole enters a steady accretion regime.
II Case study: $\ell=2$ driver
In the following, we elaborate on these phenomena by presenting additional results and interpretations for both the weak regime (Sec. II.1) and the strong regime (Sec. II.2). In particular, the discussion of the weak regime helps build intuition about our setup and draw on insights from black hole linear perturbation theory and tidal phenomena.
II.1 Weak regime
$\underbrace{(\omega,\ell)\otimes\ldots(\omega,\ell)}_{n}.$ (8)
In our case, they lead to:
  • Linear order: $\omega$ in $\ell=2$, with amplitude $\propto A_{i}$.
$|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}.$ (9)
At third order, the angular selection rules can be formulated in terms of Racah W-coefficients [37]. This structure naturally suggests that a $n$th-order nonlinear interaction driven by a $\ell=2$ mode generically populates multiple harmonics up to $\ell_{\max}=2n$. A direct consequence is that the $2\omega$ component in $\ell=6$ (see discussions below in Sec. II.3) cannot arise at quadratic order $(2\omega=\omega+\omega)$, but must instead originate from a fourth-order process: $2\omega=\omega+\omega+\omega-\omega$.
II.2 Strong regime
Figure 8 complements our discussions in the main text. Similar to the $\ell=6$ response, where $2\omega$ and $\omega$ emerge sequentially at late times, Fig. 8a shows that in $\ell=4$, the $\omega$ component is nonlinearly excited (orange region) after the laminar modes $2\omega$ and $3\omega$ (gray regions) have peaked and the driver has stabilized. As discussed in Sec. II.1 and Fig. 7, these two laminar modes arise from second- and third-order couplings, respectively. The yellow dashed curve in the bottom panel of Fig. 8a connects the dominant laminar mode $2\omega$ with the later-emerging mode $\omega$, highlighting an inverse cascade.
Within the time interval $0,700M_{i}]$, the amplitude distributions are qualitatively consistent with expectations from the normal perturbation perspective. As discussed in Sec. II.1, in $\ell=2$, each frequency multiple $n\omega$ ($n\in{\rm integer}$) arises from a $n$-th order coupling, whose amplitude decreases exponentially with $n$. In $\ell=4$ (Fig. 9b), the dominant frequency component is $2\omega$, originating at second order. The $\omega$ and $3\omega$ components both appear at third order, and thus have comparable amplitudes, together forming a “ v ” shape. Higher multiples arise from progressively higher-order interactions and show correspondingly smaller amplitudes. Finally, in $\ell=6$ (Fig. 9c), the leading third-order modes are $\omega$ and $3\omega$, followed by two fourth-order modes $2\omega$ and $4\omega$. Modes arising from the same perturbative order have comparable amplitudes, producing a “ N ” shape in the spectrum.
II.3 Injection at various frequencies
Our investigation in Sec. II.2 focused on a specific injection frequency $\omega M_{i}=0.5$, chosen somewhat arbitrarily but intentionally close to and slightly above the $\ell=2$ fundamental quasinormal mode frequency. One motivation for this choice was to ensure that a large portion of the injected wave reaches the horizon. In what follows, we demonstrate that the key features of the emergent instability persist for other low-frequency injections.
Figure 11 shows the amplitude envelopes of several modes undergoing the emergent instability, with injection frequencies $\omega M_{i}$ at 0.4 (a) and $0.6$ (b). We first observe that the $2\omega$ component in $\ell=6$ undergoes a laminar excitation during the initial injection phase ($t<t_{0}=600M_{i}$, in gray). As discussed earlier in Sec. II.1, it is a fourth-order effect: $2\omega(=\omega+\omega+\omega-\omega)$. The peak amplitude of this mode has been confirmed to scale as $A_{\rm inj}^{4}$. After the injection stabilizes $(t>t_{0}=600M_{i})$, $2\omega$ decays to some extent and undergoes the secondary excitation (open circle). The mode grows more rapidly for $\omega M_{i}=0.6$ than 0.4, and in the former case, the emergent excitation dominates over the initial laminar response. Ultimately, the evolution curve of $2\omega$ forms a two-peak structure.
III Case study: $\ell=6$ driver
III.1 Simulations
  • Linear order: $\omega$ in $\ell=6$, with amplitude $\propto A_{i}$.
III.2 A toy model
As discussed in the main text, the resonant behavior arises from a three-mode coupling. To illustrate this mechanism, consider a toy model in which a wave $\psi$ propagates on flat spacetime and is sourced by two parent modes $\psi_{1}$ and $\psi_{2}$. Then Eq. (5) in the main text becomes
$(\partial_{t}^{2}-\partial_{x}^{2})\psi=\psi_{1}\psi_{2}.$ (10)
Assume the parent modes take the form
$\psi_{1}=A_{i}^{2}e^{2i\omega(t+x)},\quad\psi_{2}=A_{i}e^{-i\omega(t+x)},$ (11)
where $\psi_{1}$ represents a quadratic mode generated by the external driver. Eq. (10) then becomes
$(\partial_{t}^{2}-\partial_{x}^{2})\psi=A_{i}^{3}e^{i\omega(t+x)}.$ (12)
Performing a Fourier transform in the $x$-direction:
$\psi(t,x)=\int\tilde{\psi}(t,k)e^{ikx}dk,$ (13)
yields
$\left(\frac{d^{2}}{dt^{2}}+k^{2}\right)\tilde{\psi}(t,k)=A_{i}^{3}e^{i\omega t}\delta(k-\omega).$ (14)
This equation admits a resonant solution
$\tilde{\psi}(t,k)=\frac{A_{i}^{3}te^{i\omega t}}{2i\omega}\delta(k-\omega).$ (15)
which, upon transforming back to real space, gives
$\psi(t,x)=\frac{A_{i}^{3}te^{i\omega(t+x)}}{2i\omega},$ (16)
The amplitude of $\psi$ therefore grows linearly in time, demonstrating the resonant nature of the three-mode coupling.
IV Simulation Methods
@@
$P_{\mu\nu}^{{\rm P}\alpha\beta}d_{t}u^{1-}_{\alpha\beta}\doteq\left({\rm incoming\penalty 10000\ gravitational\penalty 10000\ wave}\right)_{\mu\nu},$ (17)